If the length of one leg of a right triangle is three times the length of the other and the hypotenuse is 20, find the length of the shorter leg.

Answer by checkley79(3341)

A^2+B^2=C^2
A=3B
C=20
(3B)^2+B^2=C^2
9B^2=20^2
10B^2=400
B^2=400/10
B^2=40
B=SQRT40
B=6.324 ANS. FOR THE SHORTEST. LEG.
A=3*6.324
A=18.97 ANS. FOR SHORTEST LEG.
PROOF:
18.97^2+6.324^2=20^2
360+40=400
400=400